OpenStudy (anonymous):

2x – 3(x – 5 + x) = 2x – 3 Prove: x = 3 Mathematical Statement Justification 2x – 3(x - 5 + x) = 2x - 3 Given 2x – 3(x + x – 5) = 2x – 3 Commutative Property of Addition 2x – 3(2x – 5) = 2x – 3 Combine Like Terms 2x – 6x + 15 = 2x – 3 (2x – 6x) + 15 = 2x – 3 – 4x + 15 = 2x – 3 – 6x + 15 = – 3 Subtraction Property of Equality – 6x = – 18 Subtraction Property of Equality x = 3 Division Property of Equality What are the missing justification steps, in order

OpenStudy (anonymous):

From 2x - 3(2x - 5) = 2x - 3 To 2x - 6x + 15 = 2x - 3 What was done?

OpenStudy (anonymous):

the steps are Distributive Property, Associative Property of Addition, Combine Like Terms Subtraction Property of Equality, Associative Property of Addition, Addition Property of Equality Subtraction Property of Equality, Commutative Property of Addition, Addition Property of Equality Distributive Property, Commutative Property of Addition, Combine Like Terms

OpenStudy (anonymous):

i think its subtraction ,associative property, addition property of equality

OpenStudy (anonymous):

Are you guessing? ;)

OpenStudy (anonymous):

nope i just dont know if putting 2x-6x in paranthesis (2x-6x) counts as a distributive property ??

OpenStudy (anonymous):

Well, my first question: From 2x - 3(2x - 5) = 2x - 3 To 2x - 6x + 15 = 2x - 3 What was done?

OpenStudy (anonymous):

distribute you multiplied 3*2x then 3*5

OpenStudy (anonymous):

Distribute, precisely :D So in your choices, only two of them say "distributive" first.

OpenStudy (anonymous):

So the first step is "Distributive Property..." Now, From 2x - 6x + 15 = 2x - 3 To (2x – 6x) + 15 = 2x – 3 What was done?

OpenStudy (anonymous):

associative property of addiditon right?

OpenStudy (anonymous):

Associative property: Changing the grouping of the addends doesn't change the sum :D So yeah

OpenStudy (anonymous):

oh okayy thankss i didnt understand that very well but thank you i understand now soo much!1

OpenStudy (anonymous):

No problem :)